Please write the following with steps
Draw a graph of each function by first writing an equivalent exponential form and then making a table of values
I am only having trouble with writing an equivalent exponential form
y = log2x

Answer by stanbon(48569)

y = log2(x plus 1)
--------------------
Exponent is y
base is 2
inverse log is x+1
x+1 = 2^y
=============
Cheers,
Stan H.